OpenStudy (destinyyyy):

A car rental agency charges $150 per week plus $0.25 per mile to rent a car. The total cost, C, for the renting the car for one week and driving it x miles can be modeled by the formula . How many miles can you travel in one week for $175?

OpenStudy (destinyyyy):

I got 193.75 but its wrong..

OpenStudy (destinyyyy):

I plugged 175 in for x.. Was that wrong?

OpenStudy (defiance):

Hi.

OpenStudy (defiance):

First, lets only use our vial information.

OpenStudy (destinyyyy):

What? Vial?

OpenStudy (defiance):

We have $175. We automatically lose $150. Cause we rented for 1 week. We have $25 to drive with. Every mile we drive, it costs $0.25

OpenStudy (defiance):

Vial as important.

OpenStudy (destinyyyy):

Um okay.

OpenStudy (defiance):

Hold on. \[$175 - $150 = $25\] \[$25 \div $0.25 = 100\]

OpenStudy (defiance):

We can drive 100 miles in 1 week.

OpenStudy (destinyyyy):

Okay. So I should have ignored the formula it gave?

OpenStudy (defiance):

Whats the formula?

OpenStudy (destinyyyy):

Oh gosh it didnt copy it... A car rental agency charges $150 per week plus $0.25 per mile to rent a car. The total cost, C, for the renting the car for one week and driving it x miles can be modeled by the formula C=0.25x+150 . How many miles can you travel in one week for $175?

OpenStudy (defiance):

Its basiclly what I said. Just righten in another way.

OpenStudy (defiance):

C is our total cost. Which was given to us at the end "$175".

OpenStudy (defiance):

So rewright it and make it look like an equation 0.25x + 150 = 175

OpenStudy (defiance):

Since we already solved it we know x = 100.

OpenStudy (defiance):

0.25(100) + 150 = 175

OpenStudy (defiance):

Which makes 25 + 150 = 175

OpenStudy (defiance):

x is still 100. Either way you solve it.

OpenStudy (destinyyyy):

Okay I see where i messed up. Thanks.

OpenStudy (defiance):

You're welcome!